During the ~3.6-nmi transit (in dynamic positioning mode) to Site 1114 (proposed site ACE- 14A), an RCB BHA was made up with a new bit and a mechanical bit release. The ship arrived on location at 2015 hr on 12 July and a positioning beacon was deployed. However, this beacon could not be used for positioning because it had shifted into half repetition rate mode, possibly because of interference from one of the precision depth recorders (PDRs) (12 or 3.5 kHz). The ship returned to the GPS coordinates, and a second beacon was deployed at 2130 hr.

SITE 1114

Hole 1114A

While lowering the pipe to determine seafloor depth, the strong current caused the knobby drilling joints to drag in the guide horn. This allowed the heave compensator to close leading the driller to incorrectly assume the seafloor had been tagged. As a result several water cores were recovered before eventually spudding Hole 1114A at 406.5 mbsl at 0130 hr on 13 July.

The RCB Cores 1R through 31R were taken from 0 to 295.4 mbsf. The intervals l42–180 mbsf and 257–276 mbsf, were characterized by rapid ROP and extremely poor recovery (0%–4%). The small amount of recovery indicated the presence of sand in these intervals. Metamorphic rocks were first recovered at ~300.0 mbsf (Core 32R) and coring continued to 352.8 mbsf (Core 37R). While coring there were several episodes when we lost acoustic contact with the positioning beacons. A third positioning beacon was deployed at 1340 hr on 14 July, but contact with this beacon was lost almost immediately. At 1642 hr, a fourth beacon was deployed ~150 m upcurrent from Hole 1114A using the zodiac. The intent was for the beacon to drift with the >1.5 nmi/hr current while sinking to the bottom and to end up closer to the actual spud location. No further positioning problems were encountered.

While preparing to retrieve Core 37R, the hole collapsed, packing off the annulus and preventing circulation. The pipe was raised 80 m above the bottom of the hole on two separate occasions before circulation could be established. We spent 6.25 hr trying to get back to bottom to resume coring without success; therefore, we decided to run wireline logs in the portion of the hole that remained open. The bit release sleeve was shifted at 316.0 mbsf; however, the bit did not release. The pipe was picked up to 304.0 mbsf while rotating the pipe with the top drive. The bit released between 304 and 316 mbsf, and the hole was displaced with 90 bbl of sepiolite mud. The pipe was raised to 80.0 mbsf, and we rigged up for logging.

The first logging run with the triple combo was able to log from 304.0 mbsf up to ~60 mbsf. The second logging run with the FMS-sonic tool was able to log up from 299.0 mbsf. Once logging was completed (0545 hr on 16 July) we ran the open-ended drill pipe back in the hole as far as possible (107.0 mbsf ) and displaced the hole with 35 bbl of 10.5-ppg bentonite gel mud. The drill string was pulled clear of the seafloor at 0625 hr, and the four positioning beacons were recovered in 1.25 hr. At 0800 hr on 16 July, the drilling assembly cleared the rig floor and we began the transit to Site 1115 (proposed site ACE- l5A). Core recovery for Hole 1114A was 43.78 m (12%).


To 180 Site 1115

To 180 Table of Contents